Elevation 27600 feet at 7 am, person reached summit elevation 29028 at 1:25 pm.
Assuming the rate of ascent to be a linear function,
a)Find rate of ascent in feet per minute

Answer by stanbon(75887)

Elevation 27600 feet at 7 am, person reached summit elevation 29028 at 1:25 pm.
Assuming the rate of ascent to be a linear function,
a)Find rate of ascent in feet per minute
---------------------
rate = rise/time
rate = (29028-27600)/(6 hr 25 min)
rate = 1428/(6*60+45)
rate = 1428/405
rate = 3.526 ft/min
=========================
Cheers,
Stan H.